# Variation of the centered maximal operator

### v1

@@ 1. Introduction
Our first main result, Theorem 1.1, supports this conjecture. We mention that there is no need to formulate separately the discrete counterpart of Conjecture $\Delta$, since by our second main result, Theorem 2.2, inequality (1.1) and its discrete analog are equivalent. Obviously, verification of either version of the conjecture, continuous or discrete, reduces to consideration of nonnegative functions. We shall prove that a certain further reduction to nonnegative simple functions is available; see Proposition 2.3. Finally, we show that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal; see Section 4.
@@
Similarly, replacing the system of real numbers $\{t_{k}\}_{0}^{K}$ by a system of integers $-\infty<A=n_{0}<n_{1}<\dots<n_{K}=B<+\infty$, such that $n_{k}-n_{k-1}>1$ for $k=1,\ldots,K$, define
$F(n)\coloneqq a{\mathds{1}_{{(-\infty,A]}}}(n)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{(n_{k-1},n_{k})}}}(n)+b{\mathds{1}_{{[B,\infty)}}}(n),$
for $n\in\mathbb{Z}\setminus\{n_{1},\ldots,n_{K-1}\}$, and set arbitrary value for $f(n_{k})$, $11\leq k\leq K-$ 1, unless $\alpha_{k}$ and $\alpha_{k+1}$ are different from 00 when we set $f(n_{k})=0$. Then
@@
Theorem 1.1 implies (1.2) or (1.3) for all indicator functions of bounded variation, $f\colon\mathbb{R}\to\{0,1\}$ or $F\colon\mathbb{Z}\to\{0,1\}$, respectively, thus strengthening [4, Theorems 1.1 and 1.3]. Notice a separation condition encoded in the considered systems $\{t_{k}\}_{0}^{K}$ and $\{n_{k}\}_{0}^{K}$, when we set $f(t_{k})=0$ or $F(n_{k})=0$ in some circumstances depending on the behavior of $f$ or $F$ around $t_{k}$ or $n_{k}$, respectively.
@@ 2. Transference
But if the left part above holds, then there is $\delta>0$ such that $Mg(t)=M_{\geq\delta}g(y)$ for $y\in(x-\delta,x)$. Since like before $|M_{\geq\delta}g(x)-M_{\geq\delta}g(y)|\leq\frac{(x-y){\rm Var}(g)}{2\delta}$ for all such $y$, the proof is concluded. ∎
@@
It is convenient to prove a version of the above for the discrete case, $\mathbb{K}=\mathbb{Z}$, and then apply the extension procedure to obtain the result for $\mathbb{K}=\mathbb{R}$.
@@
We are ready to prove Theorem 1.1. By the extension procedure it suffices to deal with $\mathbb{K}=\mathbb{R}$. Indeed, given $F\in{\rm BV}_{a}^{b}(\mathbb{Z})$ as in Theorem 1.1 with the corresponding system $\{n_{k}\}_{0}^{K}$, define $\tilde{f}\in{\rm BV}_{a}^{b}(\mathbb{R})$ with the corresponding system $\{t_{k}\}_{0}^{K}$, $t_{k}=n_{k}$, by $\tilde{f}\coloneqq\sum_{n\in\mathbb{Z}}F(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$. Using the separation condition for $F$ and modifying $\tilde{f}$ suitably on $\mathbb{Z}+\frac{1}{2}$, we adjust $\tilde{f}$ to the form of $f$ as in Theorem 1.1 so that ${\rm Var}(f)={\rm Var}(\tilde{f})={\rm Var}(F)$ and ${\rm Var}(Mf)={\rm Var}(M\tilde{f})\geq{\rm Var}(MF)$ hold. Then (1.3) for $F$ follows from (1.2) for $f$.
@@
The same inequality easily follows when $Mf$ has exactly one substantial local maximum. Assume therefore that $N\geq 2$ and $x_{1}<\dots<x_{N}$ are all substantial local maxima of $Mf$. Since $Mf\geq\frac{a+b}{2}$, $T_{I_{k}}(\pm\infty)=0$, and $Mf$ decreases on $[x_{N},\infty)$, arguing as in the proof of Lemma 3.2 gives

### v2

@@ 1. Introduction
Our first main result, Theorem 1.1, supports this conjecture. We mention that there is no need to formulate separately the discrete counterpart of Conjecture $\Delta$, since by our second main result, Theorem 2.2, inequality (1.1) and its discrete analogue are equivalent. Clearly, verification of either version of the conjecture, continuous or discrete, reduces to consideration of nonnegative functions. We shall prove that a certain further reduction to nonnegative simple functions is available; see Proposition 2.3. Finally, we show that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal; see Section 4.
@@
Similarly, replacing the system of real numbers $\{t_{k}\}_{0}^{K}$ by a system of integers $-\infty<A=n_{0}<n_{1}<\dots<n_{K}=B<+\infty$, and assuming $\alpha_{k}\alpha_{k+1}=0$, $11\leq k\leq K-$ 1, define
$F(n)\coloneqq a{\mathds{1}_{{(-\infty,A)}}}(n)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{[n_{k-1},n_{k})}}}(n)+b{\mathds{1}_{{[B,\infty)}}}(n)$
for all $n\in\mathbb{Z}$. Then
@@
Theorem 1.1 implies (1.2) or (1.3) for all indicator functions of bounded variation, $f\colon\mathbb{R}\to\{0,1\}$ or $F\colon\mathbb{Z}\to\{0,1\}$, respectively, thus strengthening [4, Theorems 1.1 and 1.3]. Notice a separation condition encoded in the considered systems $\{t_{k}\}_{0}^{K}$ and $\{n_{k}\}_{0}^{K}$, when we require either $f(t_{k})=0$, in some circumstances depending on the behavior of $f$ around $t_{k}$, or $\alpha_{k}\alpha_{k+1}=0$, respectively.
@@ 2. Transference
But if the left part above holds, then there is $\delta>0$ such that $Mg(y)=M_{\geq\delta}g(y)$ for $y\in(x-\delta,x)$. Since like before $|M_{\geq\delta}g(x)-M_{\geq\delta}g(y)|\leq\frac{(x-y){\rm Var}(g)}{2\delta}$ for all such $y$, the proof is concluded. ∎
@@
It is convenient to prove a version of the above for the discrete case, $\mathbb{K}=\mathbb{Z}$, and then apply Theorem 2.2 and the extension procedure to obtain the result for $\mathbb{K}=\mathbb{R}$.
@@
We are ready to prove Theorem 1.1. By the extension procedure it suffices to deal with $\mathbb{K}=\mathbb{R}$. Indeed, given $F\in{\rm BV}_{a}^{b}(\mathbb{Z})$ as in Theorem 1.1 with the corresponding system $\{n_{k}\}_{0}^{K}$, define $f\in{\rm BV}_{a}^{b}(\mathbb{R})$ with the corresponding system $\{t_{k}\}_{0}^{K}$, $t_{k}=n_{k}-\frac{1}{2}$, by $f\coloneqq\sum_{n\in\mathbb{Z}}F(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$. By using the separation condition $\alpha_{k}\alpha_{k+1}=0$ for $F$, we verify that $f$ has the form as in Theorem 1.1. Moreover, ${\rm Var}(f)={\rm Var}(F)$ and ${\rm Var}(Mf)\geq{\rm Var}(MF)$. Then (1.3) for $F$ follows from (1.2) for $f$.
@@
The same conclusion easily follows when $Mf$ has exactly one substantial local maximum. Assume therefore that $N\geq 2$ and $x_{1}<\dots<x_{N}$ are all substantial local maxima of $Mf$. Since $Mf\geq\frac{a+b}{2}$, $T_{I_{k}}(\pm\infty)=0$, and $Mf$ decreases on $[x_{N},\infty)$, arguing as in the proof of Lemma 3.2 gives